OpenStudy (anonymous):

What is the solution set of the following equation? -8x + x + 15 = -7x + 12 Ø {1/7} {all reals}

OpenStudy (agreene):

Ø no solutions exist, because: (-8+1)x+15=-7x+12 -7x+15=-7x+12 15=12 <== not true.